OpenStudy (anonymous):

What is the value of z in the equation z + 7.8 = 16.4? A. 8.6 B. 7.6 C. 9.6 D. 24.2

OpenStudy (anonymous):

a

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle z + 7.8 = 16.4 }\) \(\large\color{black}{ \displaystyle z + 7.8 \color{blue}{-7.8}= 16.4\color{blue}{-7.8} }\)

OpenStudy (solomonzelman):

What did I do (with both sides) ?

OpenStudy (anonymous):

8.6+7.8=16.4 medal please

OpenStudy (keygrover):

A

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

you do 16.4 - 7.8 which =s A

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

your welcome